OpenStudy (anonymous):

solve for x: x^2+3x-40=0 ? please solve

OpenStudy (zzr0ck3r):

hint: -5*8 = -40 and -5+8 = 3

OpenStudy (anonymous):

dude please solve i want to learn how to

OpenStudy (zzr0ck3r):

well what you do is find two numbers that multiply to get -40 and add to get 3 (because -40 is the last term, and 3 is the coefficient on the middle term) once you have those numbers(which I told you are -5 and 8) you do this \((x-5)(x+8)=0\)

OpenStudy (zzr0ck3r):

this will work if it is factor-able and of the form \(x^2+bx+c=0\)

OpenStudy (anonymous):

but i need to solve for x

OpenStudy (anonymous):

@walking_stick

OpenStudy (zzr0ck3r):

x-5=0 implies x=5, x+8=0 implies x=-8

OpenStudy (zzr0ck3r):

some of this shuold look familiar to you. If you have any questions please ask.

OpenStudy (anonymous):

Notcie that (x-5)(x+8)=0. This follows the form ab=0, where (x-5) is a and (x+8) is b. Notice that if ab=0, that means that either a or b must be zero for the equation to be true. Therefore, we set a and b equal to 0 and solve for x.

OpenStudy (anonymous):

And then what ?@walking_stick

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

Well lol thats about it

OpenStudy (anonymous):

Just as zzrocker said, if solve x-5=0 and x+8=0 for x and both of these are your solutions, depending on your domain.

OpenStudy (anonymous):

Great thanks dudes